OpenStudy (anonymous):

Given w = xy + yz + xz, x = s + t, y = s - t, and z = st. Using the Chain Rule, Fin d ∂z/∂s and ∂z/∂t.

OpenStudy (anonymous):

is it ∂z/∂s = (y+yz+z)(1+t)+(x+z+xz)(1-t) ∂z/∂t = (y+yz+z)(s+1)+(x+z+xz)(s-1) ?!?! Plz help

OpenStudy (anonymous):

right

OpenStudy (anonymous):

it is

OpenStudy (loser66):

I am sorry, my knowledge is not enough to help you